# Degree-Weighted DeGroot Learning

### v2

@@ 5 Discussion on the Symmetry Assumption
Before concluding, we discuss the symmetry assumption imposed on the adjacency matrix. Although both directed and undirected networks are natural in a learning setting, we focus on undirected networks because they are required for using the second largest eigenvalue in modulus (SLEM) to bound convergence speed and, more importantly, to derive our asymptotic result in this context.
Why Symmetry is Needed.
@@
What Happens When the Network is Asymmetric.
@@
$\mathbf{P}^{\prime}=\begin{pmatrix}0.3&0.2+a\\ 0.2-a&0.3\end{pmatrix},$ (10)
@@
Figure 3a plots convergence speeds as a function of the degree dependence $\alpha$. For small asymmetry $a$, the convergence speed curves are close to the symmetric benchmark ($a=0$). When $a$ is large (e.g., $a=0.2$), the behavior differs significantly: $|\lambda_{2}|$ approaches 1 for high degree dependence $\alpha$, indicating a failure of convergence.
[…]
@@
(b) Convergence of $\mathbf{T}^{t}$, $2\log\|\mathbf{T}^{t}-(\mathbf{Q}^{-1})_{,1}\mathbf{Q}_{1,}\|_{F}$.
Figure 3: Learning Speed and Convergence of $\mathbf{T}^{t}$ With Various Levels of Asymmetry, $a$.
Notes: There are two groups with group sizes $n_{1}=500$ and $n_{2}=300$; linking probabilities are $p_{11}=p_{22}=0.3$, $p_{12}=0.2+a$, $p_{21}=0.2-a$, and $a$ captures the level of asymmetry.
Key Takeaways from the Simulations.
@@
  • A slight departure from symmetry leads to similar learning outcomes, suggesting that the symmetric case is not a knife-edge condition and that our findings exhibit at least local robustness (see Figure 3 for relatively small $a$).
  • A large departure from symmetry, however, introduces substantial differences and challenges. For instance, in symmetric matrices, all eigenvalues are real, and $|\lambda_{2}|$ is strictly less than one, whereas in asymmetric matrices, eigenvalues may be complex, and $|\lambda_{2}|$ can approach one, indicating extremely slow convergence (see Figure 3 for $a=0.18$ or $0.2$).
6 Conclusion
@@
These findings have broader implications for how societies process information with opinion leaders and centralized sources. Future research could extend the analysis to dynamic weighting rules, complex networks, or strategic belief updating. Understanding these mechanisms is crucial for designing information environments that balance rapid consensus with accurate collective learning.
Appendices
Appendix A Wisdom: General Settings and Related Proofs
A.1 General Setting for the Discussion on Wisdom
Consider a sequence of societies indexed by the population size $n$, with $n_{h}$ high-degree agents, $n_{l}=n-n_{h}$ low-degree agents, and a linking probability $q(n)$ for each pair consisting of one high-degree agent and one low-degree agent. We assume the following:
@@ A.1 General Setting for the Discussion on Wisdom
    The group of high-degree agents remains bounded: $n_{h}=O(1)<\infty$, ensuring that the low-degree agents constitute an unbounded group with $n_{l}=n-n_{h}=O(n)$.
@@
    Each low-degree agent has a bounded expected degree, i.e., $d_{l}=O(1)<\infty$.
  3. 3.
    Each high-degree agent has an unbounded expected degree, i.e., $d_{h}\sim n\,q(n)\to\infty$, given that $n_{h}$ is bounded.
Notice that we do not require either the high-degree or low-degree agents to be homogeneous—there may be multiple subgroups within each category. Moreover, the linking structures within each group can be arbitrary, as long as the above assumptions hold.
Proposition 3 (Wisdom).
Given any $\alpha$, a sequence of societies as described above is wise if and only if
$\left(nq(n)\right)^{\max\{\alpha+1,2\alpha\}}=o(n).$
Proof.
We drop $(n)$ from the notations when there is no ambiguity. Let $H$ and $L$ be the sets of high- and low-degree agents, respectively. We have $|H|=n_{1}$ and $|L|=n_{2}=n-n_{1}$.
By Eq. (7) in Proposition 1, $i$’s influence $s_{i}=S_{i}/\sum_{i^{\prime}}S_{i^{\prime}}$, where $S_{i}\equiv\sum_{j}r_{ij}d_{i}^{\alpha}d_{j}^{\alpha}$, with $r_{ij}$ being the linking probability between $i$ and $j$. By assumption, $r_{hl}=q\text{ for }h\in H$ and $l\in L$.
It follows from the conditions that, $\forall h\in H$ and $l\in L$:
  • $n_{h}=O(1)$ (cond. 1) and $n_{l}=O(n)$;
  • $d_{l}-n_{h}\leq d_{l}=O(1)$ (cond. 2);
  • $d_{h}-n_{l}q=\sum_{h^{\prime}\in H}r_{ij}\leq n_{h}=O(1)$ and $d_{h}=O(n_{l}q)=O(nq)$; therefore
  • $d_{l}^{\alpha}=O(1)$ and $d_{h}=O\left((nq(n))^{\alpha}\right)$, $\forall\alpha$.
Now we are ready to characterize $S_{h}$ and $S_{l}$, as follows:
$S_{h}$ $=n_{l}qO\left((nq(n))^{\alpha}\right)+(d_{h}-n_{l}q)O\left((nq(n))^{2\alpha}\right)$
$=O\left((nq(n))^{\alpha+1}\right)+O\left((nq(n))^{2\alpha}\right)$
$=O\left((nq(n))^{\max\{\alpha+1,2\alpha\}}\right);$
$S_{l}$ $=n_{h}qO\left((nq(n))^{\alpha}\right)+(d_{l}-n_{h}q)O(1)$
$=O\left((nq(n))^{\alpha+1}\right)+O(n).$
High-degree agents’ influences:
$s_{h}=\frac{S_{h}}{\sum_{h\in H}S_{h}+\sum_{l\in L}S_{l}}=o(1)$$\iff$ $\frac{S_{h}}{nS_{l}}=o(1)$ $\iff(nq(n))^{\max\{\alpha+1,2\alpha\}}=o(n)$.
Low-degree agents’ influences:
$s_{l}=\frac{S_{l}}{n_{h}S_{h}+n_{l}S_{l}}=o(1)$ for sure.
Thus, the society is wise iff $(nq(n))^{\max\{\alpha+1,2\alpha\}}=o(n)$. ∎
A.2 Proofs Omitted in Section 4.1
Proofs of Example 2.
There is one high-degree agent with $d_{1}=n-1$ and $n-1$ low-degree agents with $d_{2}=2$.
When $\alpha=0$, the influence of the high-degree agent is given by:
$s_{1}=\frac{d_{1}}{d_{1}+(n-1)d_{2}}=\frac{n-1}{(n-1)+2(n-1)}=\frac{1}{3}.$
When $\alpha=-1$, the influence of the high-degree agent becomes:
$s_{1}=\frac{\frac{1}{d_{1}}}{\frac{1}{d_{1}}+(n-1)\left(\frac{1}{d_{2}}\left(\frac{1}{n-1}+\frac{1}{d_{2}}\right)\right)}\approx\frac{2}{n-1}\to 0,$
as $n\to\infty$.
[…]
Proof of Example 3.
In this model, there is one high-degree agent with $d_{1}=\sqrt{n-1}$ and $n-1$ low-degree agents with expected degree $d_{2}=1+\frac{1}{\sqrt{n-1}}$.
When $\alpha=0$, the influence of the high-degree agent is given by:
$s_{1}=\frac{d_{1}}{d_{1}+(n-1)d_{2}}=\frac{\sqrt{n-1}}{\sqrt{n-1}+(n-1)}\approx\frac{1}{\sqrt{n-1}}\to 0,$
as $n\to\infty$.
When $\alpha=2$, the influence of the high-degree agent is approximately:
$s_{1}\approx\frac{\sqrt{n-1}\,(\sqrt{n-1})^{2}}{\sqrt{n-1}\,(\sqrt{n-1})^{2}+(n-1)\Bigl{(}1+\sqrt{n-1}\Bigr{)}}.$
Since $(\sqrt{n-1})^{3}=(n-1)^{3/2}$ and $(n-1)(1+\sqrt{n-1})\sim(n-1)^{3/2}$ for large $n$, it follows that:
$s_{1}\to\frac{1}{2},\quad\text{as }n\to\infty.$
[…]
Proof of Proposition 2.
The result follows directly from Proposition 3. $\Box$
Appendix B Proofs for Results in Section 3
B.1 Preliminaries: Generalized Degree-Dependence Function
@@
In this appendix, we prove that our results hold under a more general setting, where $d^{\alpha}$ is replaced by $\phi(\alpha,d)$. Here, $\phi(\alpha,d)$ satisfies Property 1–Property 3.
Property 1.
@@
Property 2.
For any two degrees $d_{1}>d_{2}$, the ratio $\phi(\alpha,d_{2})/\phi(\alpha,d_{1})$ is strictly decreasing in $\alpha$:
$\frac{d}{d\alpha}\left(\frac{\phi(\alpha,d_{2})}{\phi(\alpha,d_{1})}\right)<0\,.$ (11)
@@
$\lim\limits_{\alpha\rightarrow\infty}\frac{\phi(\alpha,d_{2})}{\phi(\alpha,d_{1})}=\lim\limits_{\alpha\rightarrow-\infty}\frac{\phi(\alpha,d_{1})}{\phi(\alpha,d_{2})}=0.$ (12)
Property 3.
@@
$\limsup_{d\rightarrow\infty}\frac{\partial\phi(\alpha,d)/\partial d}{\phi(\alpha,d)/d}<\infty\,,$ (13)
and
$\limsup_{d\rightarrow\infty}\frac{\,\partial^{2}\phi(\alpha,d)/\partial d^{2}}{(\partial\phi(\alpha,d)/\partial d)/d}<\infty\,.$ (14)
@@ B.2 Proof Preliminaries
B.2.1 $\mathbf{T}$ Similar to a Symmetric Matrix
As a preliminary step for the following proofs, we first establish that $\mathbf{T}$ is similar to a symmetric matrix.
Lemma 1.
@@ B.2.1 $\mathbf{T}$ Similar to a Symmetric Matrix
$\mathbf{T}=\mathbf{D}_{1,\mathbf{A}}^{-1}\mathbf{A}\mathbf{D}_{2,\mathbf{A}}\,,$ (15)
@@
$(\mathbf{D}_{1,\mathbf{A}})_{ii}=\sum_{j}A_{ij}\,\phi(\alpha,d_{j}(\mathbf{A}))\,,$
$(\mathbf{D}_{2,\mathbf{A}})_{ii}=\phi(\alpha,d_{i}(\mathbf{A}))\,.$
@@
$\mathbf{S}=\left(\mathbf{D}_{1,\mathbf{A}}\mathbf{D}_{2,\mathbf{A}}\right)^{-1/2}.$ (16)
@@
$\mathbf{S}^{-1}\mathbf{T}\mathbf{S}=\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}.$ (17)
Since $\mathbf{A}$ is symmetric, the right-hand side is also symmetric, proving that $\mathbf{T}$ is similar to a symmetric matrix. ∎
@@
$\mathbf{T}^{*}=\mathbf{D}^{-1}_{1,\mathbf{R}}\mathbf{R}\mathbf{D}_{2,\mathbf{R}}\,,$ (18)
where $\mathbf{R}=\mathbb{E}[\mathbf{A}]$, and $\mathbf{D}_{1,\mathbf{R}}$ and $\mathbf{D}_{2,\mathbf{R}}$ are diagonal matrices defined as
$(\mathbf{D}_{1,\mathbf{R}})_{ii}=\sum_{j}R_{ij}\,\phi(\alpha,d_{j}(\mathbf{R}))\,,$
$(\mathbf{D}_{2,\mathbf{R}})_{ii}=\phi(\alpha,d_{i}(\mathbf{R}))\,.$
B.3 Proof of Proposition 1
B.3.1 Proof of the First Part of Proposition 1
@@
Lemma 2.
For a positive, stochastic matrix $\mathbf{T}$, the spectral radius $\rho(\mathbf{T})=1$ is a simple eigenvalue. Moreover, the limit $\mathbf{T}^{\infty}$ exists and is given by
@@
where $\mathbf{v}$ and $\mathbf{w}$ are the left and right eigenvectors of $\mathbf{T}$ corresponding to $\lambda_{1}$, normalized such that $\mathbf{w}^{\top}\mathbf{v}=1$.
@@
Next, we introduce another lemma that will also be used in the proof.
Lemma 3.
@@
Proof.
Since $\mathbf{v}$ is an eigenvector of $\mathbf{M}_{1}$ with eigenvalue $\lambda$, we have
$\mathbf{M}_{1}\mathbf{v}=\lambda\mathbf{v}.$
Multiplying both sides by $\mathbf{P}$ gives
$\mathbf{P}\mathbf{M}_{1}\mathbf{v}=\lambda\mathbf{P}\mathbf{v}.$
Using the similarity relation $\mathbf{M}_{2}=\mathbf{P}\mathbf{M}_{1}\mathbf{P}^{-1}$, we obtain
$\mathbf{M}_{2}(\mathbf{P}\mathbf{v})=\mathbf{P}\mathbf{M}_{1}\mathbf{P}^{-1}(\mathbf{P}\mathbf{v})=\mathbf{P}\mathbf{M}_{1}\mathbf{v}=\lambda\mathbf{P}\mathbf{v}.$
Thus, $\mathbf{P}\mathbf{v}$ is an eigenvector of $\mathbf{M}_{2}$ corresponding to $\lambda$, as desired. ∎
@@
$\mathbf{T}^{\infty}=\lim_{t\rightarrow\infty}\mathbf{T}^{t}=\lim_{t\rightarrow\infty}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{D}_{2,\mathbf{A}}^{-1/2}\left(\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\right)^{t}\mathbf{D}_{1,\mathbf{A}}^{1/2}\mathbf{D}_{2,\mathbf{A}}^{1/2}.$ (19)
By Lemma 2, we know that the limit converges:
$\lim_{t\rightarrow\infty}\left(\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\right)^{t}=\lambda_{1}\mathbf{v}_{1}\mathbf{v}_{1}^{\top},$ (20)
where $\lambda_{1}=1$ is the largest eigenvalue in magnitude, and $\mathbf{v}_{1}$ is the corresponding unit eigenvector of the matrix $\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}$. Since the matrix in (20) is symmetric, its left and right eigenvectors are identical.
To determine $\mathbf{v}_{1}$, we apply Lemma 3. It is straightforward to verify that $\mathbf{D}_{1,\mathbf{A}}^{-1}\mathbf{A}\mathbf{D}_{2,\mathbf{A}}$ has the eigenvector $\mathbf{e}=(1,\dots,1)^{\top}$ corresponding to eigenvalue $1$. Consequently, the matrix $\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}$ has the eigenvector
@@
Thus, the unit eigenvector is given by $\mathbf{v}_{1}=\mathbf{w}_{1}/\|\mathbf{w}_{1}\|_{2}$, and substituting into (19) yields
@@
$=\|\mathbf{v}_{1}\|_{2}^{-2}\mathbf{E}_{n}\mathbf{D}_{1,\mathbf{A}}\mathbf{D}_{2,\mathbf{A}},$

### v3

@@ 5 Discussion on the Symmetry Assumption
Before concluding, we discuss the symmetry assumption imposed on the adjacency matrix. Although both directed and undirected networks are natural in a learning setting, we focus on undirected networks because they are required for using the second-largest eigenvalue in modulus (SLEM) to bound convergence speed and, more importantly, to derive our asymptotic result in this context.
Why Symmetry is Needed
@@
What Happens When the Network is Asymmetric
@@
(8) $\mathbf{P}^{\prime}=\begin{pmatrix}0.3&0.2+a\\ 0.2-a&0.3\end{pmatrix}$,
@@
Figure 3(a) plots convergence speeds as a function of the degree dependence $\alpha$. For small asymmetry $a$, the convergence speed curves are close to the symmetric benchmark ($a=0$). When $a$ is large (e.g., $a=0.2$), the behavior differs significantly: $|\lambda_{2}|$ approaches 1 for high degree dependence $\alpha$, indicating a failure of convergence.
Figure 3(b) depicts the convergence of $\mathbf{T}^{t}$ and complements the above observation. In particular, we plot the logarithm of the Frobenius norm, $2\log\|\mathbf{T}^{t}-(\mathbf{Q}^{-1})_{,1}\mathbf{Q}_{1,}\|_{F}$, for $t=1,\dots,20$ at $\alpha=1$. [^29] If $\mathbf{T}$ is diagonalizable, then $\mathbf{T}=\mathbf{Q}^{-1}\Lambda\mathbf{Q}$, leading to $\mathbf{T}^{\infty}=(\mathbf{Q}^{-1})_{,1}\mathbf{Q}_{1,}$, where $\mathbf{Q}_{1,}$ is the first row of $\mathbf{Q}$ and $(\mathbf{Q}^{-1})_{,1}$ is the first column of $\mathbf{Q}^{-1}$. With mild asymmetry (say $a\leq 0.05$), $\|\mathbf{T}^{t}-(\mathbf{Q}^{-1})_{,1}\mathbf{Q}_{1,}\|_{F}<e^{-10/2},t\in[8,20]$, indicating convergence of $\mathbf{T}^{t}$ to the limit matrix. With large asymmetry ($a=0.18$ or $0.2$), $\|\mathbf{T}^{t}-(\mathbf{Q}^{-1})_{,1}\mathbf{Q}_{1,}\|_{F}>e^{0}=1,t\in[1,20]$, meaning $\mathbf{T}^{t}$ does not appear to converge.
@@
(b) Convergence of $\mathbf{T}^{t}$.
Figure 3: Learning Speed and Convergence of $\mathbf{T}^{t}$ with Varying Asymmetry, $a$.
{figurenotes}
There are two groups with group sizes $n_{1}=500$ and $n_{2}=300$; linking probabilities are $p_{11}=p_{22}=0.3$, $p_{12}=0.2+a$, $p_{21}=0.2-a$, and $a$ captures the level of asymmetry. Panel (b) plots the logarithm of the Frobenius norm, $2\log\|\mathbf{T}^{t}-(\mathbf{Q}^{-1})_{,1}\mathbf{Q}_{1,}\|_{F}$.
@@
  • A slight departure from symmetry leads to similar learning outcomes, suggesting that the symmetric case is not a knife-edge condition and that our findings exhibit at least local robustness (Figure 3 with relatively small $a$).
  • A large departure from symmetry, however, introduces substantial differences and challenges. For instance, in symmetric matrices, all eigenvalues are real, and $|\lambda_{2}|$ is strictly less than one, whereas in asymmetric matrices, eigenvalues may be complex, and $|\lambda_{2}|$ can approach one, indicating extremely slow convergence (Figure 3 with $a=0.18$ or $0.2$).
6 Concluding Remarks
@@
The economic implications of the trade-off between wisdom and convergence speed offer intriguing avenues for further exploration. For instance, future work could investigate scenarios in which policymakers or platform designers prioritize rapid consensus formation at the expense of societal wisdom. It would also be valuable to characterize optimal threshold values of the degree-dependence in specific informational or social contexts, thereby bridging theoretical predictions with practical applications.
Appendix A Appendix A. Primitivity of $\mathbf{R}(n)$, $\mathbf{T}^{*}(n)$ and $\mathbf{A}(n)$
Proposition A.1
Under Assumptions 13, there exists an integer $n_{0}\in\mathbb{N}$ s.t. $\forall n\geq n_{0}$:
@@
    The expected adjacency matrix $\mathbf{R}(n)$ and learning matrix $\mathbf{T}^{*}(n)$ are primitive.
@@
    The realized adjacency matrix $\mathbf{A}(n)$ is primitive with probability at least $1-16n^{-2}$.
Proof of Proposition A.1
First, we establish the primitivity of the expected matrices $\mathbf{T}^{*}$ and $\mathbf{R}$. Assumption 3 states that $\limsup_{n\rightarrow\infty}|\lambda_{2}(\mathbf{T}^{*})|<1$. By the definition of limit superior, there exists a constant $c<1$ and an integer $n_{1}$ such that for all $n\geq n_{1}$, we have $|\lambda_{2}(\mathbf{T}^{*}(n))|\leq c<1$. Since $\mathbf{T}^{*}$ is a row-stochastic matrix, its largest eigenvalue is 1. The condition $|\lambda_{2}(\mathbf{T}^{*})|<1$ implies that the eigenvalue 1 is simple and strictly dominant in modulus. For a nonnegative matrix, this condition is sufficient to guarantee primitivity (irreducibility and aperiodicity). Thus, $\mathbf{T}^{*}(n)$ is primitive for all $n\geq n_{1}$. Since $\mathbf{T}^{*}$ and $\mathbf{R}$ share the same zero-pattern (connectivity structure) by construction, the primitivity of $\mathbf{T}^{*}$ implies that $\mathbf{R}$ is also primitive for all $n\geq n_{1}$.
Next, we consider the realized adjacency matrix $\mathbf{A}$ and learning matrix $\mathbf{T}$. By the eigenvalue concentration results established in Part 2 of Theorem 1, [^30] We remark that the proof of Part 2 of Theorem 1 does not rely on the realized matrices $\mathbf{A}$ and $\mathbf{T}$ being primitive with a high probability. Therefore, using that result in the present proof does not involve circular reasoning. the second-largest eigenvalue of the realized learning matrix, $|\lambda_{2}(\mathbf{T})|$, converges in probability to $|\lambda_{2}(\mathbf{T}^{*})|$. $\forall n\geq n_{1}$ (defined in the previous paragraph), since $|\lambda_{2}(\mathbf{T}^{*})|\leq c<1$, the spectral gap $1-|\lambda_{2}(\mathbf{T}^{*})|\geq 1-c>0$ is bounded away from zero. Consequently, there exists an integer $n_{0}\geq n_{1}$ such that for all $n\geq n_{0}$, the event
$E_{n}=\{|\lambda_{2}(\mathbf{T})|<1\}$
occurs with probability at least $1-16n^{-2}$. Conditioning on the event $E_{n}$, since $\mathbf{T}$ is row-stochastic, the strict inequality $|\lambda_{2}(\mathbf{T})|<1$ implies that $\mathbf{T}$ is primitive. Finally, since each realized $\mathbf{A}$ and its corresponding $\mathbf{T}$ share the exact same support (i.e., $\mathbf{T}_{ij}>0$ if and only if $\mathbf{A}_{ij}>0$), $\mathbf{T}$ is primitive if and only if $\mathbf{A}$ is primitive. It follows that for all $n\geq n_{0}$:
$\mathbb{P}(\mathbf{A}\text{ is primitive})=\mathbb{P}(\mathbf{T}\text{ is primitive})\geq\mathbb{P}(E_{n})\geq 1-16n^{-2}.$
This completes the proof. $\Box$
Appendix B Appendix B. Proofs of Results in Section 3
1 Preliminaries: Generalized Degree-Dependence Function
@@
In this section, we prove that our results hold under a more general setting, where $d^{\alpha}$ is replaced by $\phi(\alpha,d)$. Here, $\phi(\alpha,d)$ satisfies Properties 1-3.
Property 1
@@
Property 2
For any $d_{1}>d_{2}$, the ratio $\phi(\alpha,d_{2})/\phi(\alpha,d_{1})$ is strictly decreasing in $\alpha$:
(B.1) $\frac{d}{d\alpha}\left(\frac{\phi(\alpha,d_{2})}{\phi(\alpha,d_{1})}\right)<0\,.$
@@
(B.2) $\lim\limits_{\alpha\rightarrow\infty}\frac{\phi(\alpha,d_{2})}{\phi(\alpha,d_{1})}=\lim\limits_{\alpha\rightarrow-\infty}\frac{\phi(\alpha,d_{1})}{\phi(\alpha,d_{2})}=0.$
Property 3
@@
(B.3) $\limsup_{d\rightarrow\infty}\frac{\partial\phi(\alpha,d)/\partial d}{\phi(\alpha,d)/d}<\infty\,$,
(B.4) $\limsup_{d\rightarrow\infty}\frac{\,\partial^{2}\phi(\alpha,d)/\partial d^{2}}{(\partial\phi(\alpha,d)/\partial d)/d}<\infty\,.$
@@ 2 Proof Preliminaries
2.1 $\mathbf{T}$ Similar to a Symmetric Matrix
As a preliminary step, we first establish that $\mathbf{T}$ is similar to a symmetric matrix.
Lemma B.1
@@ 2.1 $\mathbf{T}$ Similar to a Symmetric Matrix
(B.5) $\mathbf{T}=\mathbf{D}_{1,\mathbf{A}}^{-1}\mathbf{A}\mathbf{D}_{2,\mathbf{A}}\,$,
@@
$(\mathbf{D}_{1,\mathbf{A}})_{ii}=\sum_{j}A_{ij}\,\phi(\alpha,d_{j}(\mathbf{A}))\,,\ \ \ (\mathbf{D}_{2,\mathbf{A}})_{ii}=\phi(\alpha,d_{i}(\mathbf{A}))\,.$
@@
(B.6) $\mathbf{S}=\left(\mathbf{D}_{1,\mathbf{A}}\mathbf{D}_{2,\mathbf{A}}\right)^{-1/2}.$
@@
(B.7) $\mathbf{S}^{-1}\mathbf{T}\mathbf{S}=\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}.$
Since $\mathbf{A}$ is symmetric, the right-hand side is also symmetric, proving that $\mathbf{T}$ is similar to a symmetric matrix. $\Box$
@@
(B.8) $\mathbf{T}^{*}=\mathbf{D}^{-1}_{1,\mathbf{R}}\mathbf{R}\mathbf{D}_{2,\mathbf{R}}\,$,
where $\mathbf{R}={\rm I}\kern-1.79993pt{\rm E}\mathbf{A}$, and $\mathbf{D}_{1,\mathbf{R}}$ and $\mathbf{D}_{2,\mathbf{R}}$ are diagonal matrices defined as
$(\mathbf{D}_{1,\mathbf{R}})_{ii}=\sum_{j}R_{ij}\,\phi(\alpha,d_{j}(\mathbf{R}))\,,\ \ \ (\mathbf{D}_{2,\mathbf{R}})_{ii}=\phi(\alpha,d_{i}(\mathbf{R}))\,.$
3 Proof of Proposition 1
3.1 Proof of the First Part of Proposition 1
@@
Lemma B.2
Let $\mathbf{T}$ be a primitive matrix with nonnegative entries. Then the spectral radius $\rho(\mathbf{T})=1$ is a simple eigenvalue. Moreover, the limit $\mathbf{T}^{\infty}$ exists and is given by
@@
where $\mathbf{v}$ and $\mathbf{w}$ are the left and right unit eigenvectors of $\mathbf{T}$ corresponding to $\lambda_{1}$.
@@
Lemma B.3
@@
Proof of Lemma B.3
Since $\mathbf{v}$ is an eigenvector of $\mathbf{M}_{1}$ with eigenvalue $\lambda$, we have $\mathbf{M}_{1}\mathbf{v}=\lambda\mathbf{v}.$ Multiplying both sides by $\mathbf{P}$ gives $\mathbf{P}\mathbf{M}_{1}\mathbf{v}=\lambda\mathbf{P}\mathbf{v}.$ Using the similarity relation $\mathbf{M}_{2}=\mathbf{P}\mathbf{M}_{1}\mathbf{P}^{-1}$, we obtain $\mathbf{M}_{2}(\mathbf{P}\mathbf{v})=\mathbf{P}\mathbf{M}_{1}\mathbf{P}^{-1}(\mathbf{P}\mathbf{v})=\mathbf{P}\mathbf{M}_{1}\mathbf{v}=\lambda\mathbf{P}\mathbf{v}.$
Thus, $\mathbf{P}\mathbf{v}$ is an eigenvector of $\mathbf{M}_{2}$ corresponding to $\lambda$, as desired. $\Box$
@@
(B.9) $\mathbf{T}^{\infty}=\lim_{t\rightarrow\infty}\mathbf{T}^{t}=\lim_{t\rightarrow\infty}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{D}_{2,\mathbf{A}}^{-1/2}\left(\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\right)^{t}\mathbf{D}_{1,\mathbf{A}}^{1/2}\mathbf{D}_{2,\mathbf{A}}^{1/2}.$
By Lemma B.2, we know that the limit converges:
(B.10) $\lim_{t\rightarrow\infty}\left(\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\right)^{t}=\lambda_{1}\mathbf{v}_{1}\mathbf{v}_{1}^{\top},$
where $\lambda_{1}=1$ is the largest eigenvalue in magnitude, and $\mathbf{v}_{1}$ is the corresponding unit eigenvector of the matrix $\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}$. Since the matrix in (B.10) is symmetric, its left and right eigenvectors are identical.
To determine $\mathbf{v}_{1}$, we apply Lemma B.3. It is straightforward to verify that $\mathbf{D}_{1,\mathbf{A}}^{-1}\mathbf{A}\mathbf{D}_{2,\mathbf{A}}$ has the eigenvector $\mathbf{e}=(1,\dots,1)^{\top}$ corresponding to eigenvalue $1$. Consequently, the matrix $\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}$ has the eigenvector
@@
Thus, the unit eigenvector is given by $\mathbf{v}_{1}=\mathbf{w}_{1}/\|\mathbf{w}_{1}\|_{2}$, and substituting into (B.9) yields
@@
$=\|\mathbf{w}_{1}\|_{2}^{-2}\mathbf{E}_{n}\mathbf{D}_{1,\mathbf{A}}\mathbf{D}_{2,\mathbf{A}},$